OpenStudy (howard-wolowitz):

Can someone help me here:

OpenStudy (michele_laino):

please be polite with moderators, otherwise I don't help you! @Howard-Wolowitz

OpenStudy (radar):

The problem seems to leave the amount of time up to the person making draft and further seems to indicate that the period of time is the same for each graph.

OpenStudy (radar):

*draft - graph

OpenStudy (michele_laino):

Let's suppose be N the number of vehicle in 2° and central ave and let's suppose that T is time within we have counted that number N of vehicles Then number of vehicle per unit timein 2° and central Ave is: N/T

OpenStudy (michele_laino):

namely N/T is the traffic flows in 2° and Central Ave

OpenStudy (michele_laino):

From the text of your problem we can write that the traffic flow in 1st Ave and High St, is twice with respect to the traffic flow in 2nd and central Ave, so we can write: traffic flows in 1st and High street is: 2(N/T) Now we have: \[\Large 2\frac{N}{T} = \frac{N}{{\left( {T/2} \right)}}\]

OpenStudy (howard-wolowitz):

alright I gotcha so far

OpenStudy (howard-wolowitz):

but what if the streets over-lapped

OpenStudy (michele_laino):

I don't know, in this case our streets are different streets

OpenStudy (michele_laino):

so, what is the right option?

OpenStudy (howard-wolowitz):

Well B

OpenStudy (michele_laino):

are you sure? reassuming: the traffic flow in 2nd and central Ave, is: \[\Large\frac{N}{T}\] whereas, the traffic flow at 1st and High street is: \[\Large \frac{N}{{\left( {T/2} \right)}}\] please look at the denominators of both fractions

OpenStudy (howard-wolowitz):

So it would be half as long?

OpenStudy (michele_laino):

yes! So what is the right option, please?

OpenStudy (howard-wolowitz):

A

OpenStudy (michele_laino):

that's right!

OpenStudy (howard-wolowitz):

awesome thanks for the help

OpenStudy (michele_laino):

thanks! :)